Circular No. 7984

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN NGC 205 S. Nakano, Sumoto, Japan, reports the discovery of a possible nova in

NGC 205 by K. Itagaki (Teppo-cho, Yamagata, Japan) on seven unfiltered CCD frames taken around Oct. 5.56 UT with a 0.60-m f/6.0 reflector. Itagaki reports that the object brightened from mag 16.4 to 16.1 in ∼ 1 hr, and he provided the following approximate position: α = 0^{h}40^{m}15^{s}.8, δ = +41^{o}44′20′′ (equinox 2000.0), which is ∼ 3′35′′ from the center of NGC 205 (= M110) in p.a. 330^{o}. A confirming unfiltered CCD image was obtained by R. Kushida with a 0.40-m reflector at Yatsugatake South Base Observatory on Oct. 5.764, showing the new object at mag 15.8 and at precise-position end figures 15^{s}.26, 19′′.8 (measured by Y. Kushida), which is 77′′ west and 193′′ north of the center of NGC 205. Nothing appears at this position to mag 18 on survey frames taken by Itagaki on May 28 and July 30 and to mag 19 on Kushida’s image taken on Oct. 3.65.

COMETS C/2002 R4, R5, R6, R7, R8, S3, S4, S5 (SOHO) Further to IAUC 7975, D. Hammer reports his measurements for sev-

eral additional comets found on SOHO website images (C3 coronagraph, except where noted) by R. Kracht (C/2002 R4, C/2002 R8, C/2002 S5; via the C2 coronagraph only), R. Matson (C/2002 R5 and C/2002 S4, the latter also visible in C2), D. Evans (C/2002 R6), J. Sachs (C/2002 R7), and X. Leprette and S. Hönig (C/2002 S3). The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below. C/2002 R4 belongs to the Marsden group (cf. IAUC 7832), and C/2002 R8, though weakly determined, evidently belongs to the Meyer group; C/2002 R5 appears to be related to C/1999 R1, a point first made by Kracht; C/2002 S4 and C/2002 S5 are members of the Kracht family (IAUC 7952). The rest are Kreutz sungrazers.

Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 R4 Sept. 3.483  10^{h}43.^{m}1  + 9^{o}07′  2002-S35
C/2002 R5  5.204  10 50.4  + 7 05  2002-S35
C/2002 R6  13.596  11 10.3  + 3 12  2002-S35
C/2002 R7  13.571  11 04.8  + 3 08  2002-S35
C/2002 R8  15.629  11 27.9  + 4 51  2002-S35
C/2002 S3  17.363  11 17.7  + 1 34  2002-S36
C/2002 S4  17.613  11 28.4  + 3 14  2002-T16
C/2002 S5  17.646  11 41.5  + 2 48  2002-T16
2002 October 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green